OpenStudy (anonymous):

The function p(x) = 2x +1 determines how many pizzas need to be purchased for an after school meeting, where x is the number of students at the meeting. The club leader uses r(p(x)) to find the amount of money to bring for the pizza purchase. The function r(x) = 4x − 3. Solve for how much money to bring when there are 9 students in the meeting. 19 33 73 84

OpenStudy (anonymous):

73 right ? @texaschic101

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Nnesha

OpenStudy (phi):

yes, 73 x is the number of students 9 students p(x) = 2x +1 and p(9)= 2*9+1 = 19 r(p(x)) to find the amount of money where r(x)= 4x − 3 r(19) = 4*19-3 = 73